A random sample of 14 cigarettes of a certain brand has an average nicotine content of 4.4 milligrams and a standard deviation of 1.9 milligrams. What is the margin of error of the 99% confidence interval for the average nicotine content of the cigarettes.



1
Expert's answer
2022-02-24T12:38:45-0500

The critical value for "\alpha = 0.01" and "df = n-1 = 13" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} =3.012275."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

The margin of error is


"Margin \ of\ error =t_c\times\dfrac{s}{\sqrt{n}}"

"=3.012275\times\dfrac{1.9}{\sqrt{14}}\approx1.5296"

The margin of error of the 99% confidence interval for the average nicotine content of the cigarettes is 1.5296 milligrams.